OpenStudy (lazyboy):

What is the function that describes the following function table? X Y 1 2 2 3 3 4 Question 5 answers The output value Y is one less than the input value X. The output value Y is two less than the input value X. The output value Y is one more than the input value X. The input value X is twice as much as the output value Y.

OpenStudy (lazyboy):

It is a table again so: x = y 1 = 2 2 = 3 3 = 4

terenzreignz (terenzreignz):

Which of the four choices is true about the inputs x and the outputs y?

OpenStudy (lazyboy):

It should be C

terenzreignz (terenzreignz):

There you go. You've answered yourself again :P

OpenStudy (lazyboy):

I know most of the answers but I like feedback :D

terenzreignz (terenzreignz):

Okay.

OpenStudy (lazyboy):

yeah

OpenStudy (lazyboy):

let me just put the next question in here ok?

OpenStudy (lazyboy):

@terenzreignz

terenzreignz (terenzreignz):

Okay.

OpenStudy (lazyboy):

When Y = 50, what is X in the function, sqrt{?} 100x - 50 = y?

terenzreignz (terenzreignz):

\[\large \sqrt{100x-50}=y\] ?

OpenStudy (lazyboy):

\[\sqrt{100}\] x - 50 = y

OpenStudy (lazyboy):

yeah

terenzreignz (terenzreignz):

Okay, we have a value for y, it's 50... \[\large \sqrt{100x-50}=50\] Now solve for x :P

OpenStudy (lazyboy):

yeah I know but I don't know what X can possibly be

terenzreignz (terenzreignz):

Find out. First, get rid of that annoying square root sign, by squaring both sides.

OpenStudy (lazyboy):

hmm.. so A?

terenzreignz (terenzreignz):

How should I know? You didn't tell me the choices.

OpenStudy (lazyboy):

uups true sry

OpenStudy (lazyboy):

0 10 25.5 100

terenzreignz (terenzreignz):

Solve for x, when you square both sides, you get \[\huge \left(\sqrt{100x-50}\right)^2=50^2\]

OpenStudy (lazyboy):

I think it is B

terenzreignz (terenzreignz):

Could you show how you arrived at that answer?

OpenStudy (lazyboy):

by guessing .____. I am not good at this :/

OpenStudy (lazyboy):

dang and this is my last questioon :o

terenzreignz (terenzreignz):

Well, remember that the square root and the square (exponent 2) cancel out, so you get \[\huge 100x-50=2500\] Now solve for x.

OpenStudy (lazyboy):

25.5 so C thanks bro :)

terenzreignz (terenzreignz):

Okay, no problem

OpenStudy (lazyboy):

NOO it was 10

OpenStudy (lazyboy):

but thanks anyway everything else was right :)

terenzreignz (terenzreignz):

It should be 25.5

OpenStudy (lazyboy):

no, the answers said 10 but no big deal

terenzreignz (terenzreignz):

Unless I was working on the wrong problem.

terenzreignz (terenzreignz):

did it really look like this \[\sqrt{100x-50}=y\]

OpenStudy (lazyboy):

yeah

terenzreignz (terenzreignz):

Well, it should be 25.5. Now, if it were this \[\huge \sqrt{100}x-50=y\] Then it would make sense that the answer is 10

OpenStudy (lazyboy):

ups yeah I thought there would be no difference my fault lol

terenzreignz (terenzreignz):

Oh well...

OpenStudy (lazyboy):

yeah anyway thanks a lot :D